OpenStudy (anonymous):

Graph the function y=2cos(x+3pi)+1

OpenStudy (anonymous):

no luck?

OpenStudy (anonymous):

here might be an easier one that i need also,two coterminal angles, a positive and a negative for theta =3pi/4

OpenStudy (anonymous):

y=2cos(x+3pi)+1 y=2 (cos(x)+cos(3pi)) +1 y=2 cos(x)+2cos(3pi) +1 y=2 cos(x)+2(-1) +1 y=2 cos(x)+-2 +1 y = 1-2cos(x) <- simplified equation

OpenStudy (anonymous):

so to graph it, just start plugging in some coordinates for x and see what you get.. so plug in: x=0, thus y=-1 x=pi, thus y=3 x=2pi, thus y=-1 ya?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

since you are working with a trig funtion (cos) that is why i say use x as pi, and not a number like 1 or 2, fyi cheers!

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

you bet!